### v1

@@ 7 Discussion
Practical Use for ML Model Developers. For developers of RL-based cyber agents, this framework serves as a diagnostic tool for understanding training behaviour. It enables:
  • Visualisation of Q-value trajectories to trace emerging action preferences,
  • Identification of key transitions via TD-error spikes and PER prioritisation,
  • Detection of brittle convergence or excessive determinism in policies,
  • Evaluating exploration strategies by behavioural richness over performance.
These insights are especially valuable in partially observable settings like CyberBattleSim, where agent performance may hinge on sparse rewards and incomplete visibility. Our framework supports more interpretable, resilient policy development by shedding light on not just what the agent learns, but how and when.
@@
Limitations and Future Work. While CyberBattleSim offers a controlled and extensible environment, it simplifies many real-world complexities, such as concurrent adversaries, deception-aware agents, and sophisticated evasion tactics. Moreover, CyberBattleSim’s network architecture, while valuable for controlled experimentation, is relatively simple and lacks the heterogeneity and complexity of real-world enterprise networks, such as layered access controls, segmented subnets, and diverse user behaviours. Additionally, our analysis currently focuses on a single-agent attacker. Future work will extend this framework to multi-agent scenarios, enabling analysis of competitive or cooperative dynamics between attackers and defenders. We also plan to conduct user studies with security analysts and red-teamers to refine how explanations are delivered, ensuring insights are both technically sound and practically usable.
8 Conclusion
@@
In this paper, we introduced a unified, multi-layer explainability framework for demystifying the decision-making of RL-based attacker agents in cyber environments. By combining MDP- and policy-level analysis, the framework offers fine-grained, temporally grounded insights into how adversaries explore networks, adapt, and exploit network vulnerabilities. Through experiments in CyberBattleSim, we demonstrated how the framework reveals strategy shifts, learning patterns, and decision inflection points often hidden in standard RL evaluations. Unlike prior post-hoc or domain-specific work, our agent-agnostic approach generalises to defender agents, enabling unified interpretation to both offensive and defensive strategies. It equips red teams with visibility into how RL agents develop adversarial tactics and offers blue teams a powerful tool to interpret and counter evolving threats. This work lays the foundation for transparent, trustworthy, and operationally useful RL in cybersecurity. Future directions include real-time integration, multi-agent extensions, and application in high-stakes AI safety contexts.

### v2

@@ VII Discussion
Practical Use for ML Model Developers. The primary beneficiaries of this framework are developers of RL-based autonomous attacker agents. The framework serves as a training-time diagnostic tool for understanding, debugging, and validating learning behaviour. It enables:
  • Visualisation of Q-value trajectories to trace how action preferences emerge, stabilise, or shift during training, distinguishing genuine strategy formation from early bias.
  • Identification of critical learning transitions via TD-error spikes and PER prioritisation, highlighting experiences that disproportionately influence policy updates.
  • Detection of brittle convergence or excessive determinism through internal confidence and entropy signals, exposing premature policy collapse even when rewards appear stable.
  • Evaluation of exploration strategies based on behavioural richness and adaptability rather than cumulative reward alone.
These capabilities are especially valuable in partially observable environments, where incomplete visibility make reward curves a poor proxy for learning health. By exposing internal learning dynamics, the framework supports the development of more interpretable and resilient attacker models, clarifying not only what the agent learns, but how and when that learning occurs.
Implications for Defensive Simulation. While our framework is primarily intended for developers of autonomous attacker agents, we emphasise that its core contributions lie in attacker modelling and training-time diagnosis. Nevertheless, the same diagnostic signals may provide useful context for defensive simulation during offline training and scenario development. As agentic attackers become more sophisticated, understanding not only what strategies emerge but also how and when they adapt can support more informed threat modelling and defensive preparedness.
Example: In a simulated environment such as CC500, defenders may observe recurring spikes in transition priority at specific stages of training, indicating moments where the attacker model shifts from reconnaissance-driven behaviour to credential reuse or lateral movement. Although such internal learning signals are not observable in live attacks, they can inform defensive simulation by highlighting which phases of an attack lifecycle are most sensitive to adaptation. This insight can be used to design targeted defensive stress tests, evaluate the robustness of segmentation or access-control policies, and assess how defensive assumptions hold under adaptive attacker behaviour.
@@
Limitations and Future Work. While CyberBattleSim provides a controlled and extensible environment for systematic experimentation, it abstracts several real-world complexities. In particular, it does not fully capture concurrent adversaries, deception-aware agents, or sophisticated evasion tactics, and its network architectures lack the heterogeneity of operational enterprise systems, such as layered access controls, segmented subnets, and diverse user behaviours. These simplifications are necessary for controlled analysis but limit direct extrapolation to real-world deployments. Moreover, our current analysis focuses on a single autonomous attacker operating in isolation. Future work will extend the framework to multi-agent settings to investigate co-evolving attacker–defender dynamics and emergent behaviours under strategic interaction. We also plan to conduct user studies with security analysts to refine how explanations are presented and interpreted, ensuring that the resulting insights are both technically grounded and practically usable.
Key Takeaways. Understanding autonomous cyber attackers requires visibility beyond final performance metrics. Our analysis shows that reward curves alone provide an incomplete, and in some cases misleading, view of learning behaviour in adversarial cyber environments. Effective evaluation therefore requires insight into learning dynamics at multiple levels of abstraction. In particular, strategic signals reveal phase transitions from reconnaissance to exploitation as environments evolve under uncertainty, while policy-level signals expose how action preferences emerge, stabilise, or collapse during training. Together, these perspectives clarify not only what strategies are learned, but critically how and when they form.
Across large-scale environments, we find that learning failure does not originate in declining reward, but emerges internally during training as exploration capacity progressively contracts. As policy entropy decreases and decision confidence sharpens, agents become locked into narrow, brittle behaviours, even while task-level rewards remain stable or improve. This decoupling demonstrates that reward curves can actively mask underlying learning pathologies rather than reveal them.
Within this context, our results expose a concrete replay-induced failure mode. Contrary to common assumptions, Prioritised Experience Replay does not reliably accelerate learning or promote earlier lateral movement in sparse, partially observable cyber environments. Instead, it oversamples transient, high-TD-error transitions, introducing replay bias that distorts value propagation and destabilises convergence in large environments. These effects are detectable only through training-time diagnostic signals, well before they manifest as observable performance collapse.
Taken together, these findings suggest that evaluating autonomous cyber agents solely on final performance is insufficient. Training-time explainability must be treated as a first-class diagnostic tool when developing, benchmarking, and validating reinforcement learning–based attacker models, particularly in sparse, partially observable cybersecurity settings.
VIII Conclusion
@@
In this paper, we introduced a unified, multi-layer explainability framework for demystifying the decision-making of RL-based attacker agents in cyber environments. By combining MDP- and policy-level analysis, the framework offers fine-grained, temporally grounded insights into how adversaries explore networks, adapt, and exploit network vulnerabilities. Through experiments in CyberBattleSim, we demonstrated how the framework reveals strategy shifts, learning patterns, and decision inflection points often hidden in standard RL evaluations. Unlike prior post-hoc only or domain-specific work, our agent-agnostic approach generalises to defender agents, enabling unified interpretation of both offensive and defensive strategies. It equips red teams with visibility into how RL agents develop adversarial tactics and offers blue teams a powerful tool to interpret and counter evolving threats. This work lays the foundation for transparent, trustworthy, and operationally useful RL in cybersecurity. While focused on adversarial cyber operations, our framework is readily applicable to other sequential decision-making domains, such as autonomous drone navigation, financial fraud detection, and critical infrastructure protection, where explainability and behavioural traceability are crucial. Future directions include real-time integration, multi-agent extensions, and application in high-stakes AI safety contexts.
Appendix
TABLE IV: Action indices and descriptions for CTF environment. These indices (1-18) correspond to the x-axis in Figure 6 and are used to interpret the emergence of action preferences.
Index Abstract Action Description
1 local(ScanBashHistory) Scan bash history for host references
2 local(ScanExplorerRecentFiles) Scan Windows recent-files list
3 local(SudoAttempt) Attempt sudo privilege escalation
4 local(ExfiltrateFlag) Read local flag file
5 local(CrackKeePassX) Crack KeepPassX vault (Linux)
6 remote(ProbeLinux) Probe to confirm Linux OS
7 remote(ProbeWindows) Probe to confirm Windows OS
8 remote(ProbeSQLServer) Probe for SQL-Server service
9 connect(HTTPS) Connect to HTTPS service
10 connect(GIT) Connect to Git service
11 connect(SSH) SSH connection with cached creds
12 connect(RDP) RDP connection with cached creds
13 connect(MySQL) Connect to MySQL service
14 connect(SSH-key) SSH using leaked key
15 connect(PING) ICMP ping
16 connect(su) Command/connection via su creds
17 remote(ProbeFlagServer) Probe the flag server
18 local(CrackKeePass) Crack KeepPass vault (Windows)
TABLE V: Agent state at Episode 20, Step 956, with five owned nodes and no discovered nodes, reflecting an early stage of exploration under partial observability.
id status properties local_attacks remote_attacks
client owned [] [SearchEdgeHistory] []
Website owned [MySql, Ubuntu, nginx/1.10.3] [CredScanBashHistory] [ScanPageSource, ScanPageContent]
Website[user=monitor] owned [MySql, Ubuntu, nginx/1.10.3] [CredScan-HomeDirectory] []
AzureStorage owned [CTFFLAG:LeakedCustomerData] [] [AccessDataWithSASToken]
AzureResourceManager owned [CTFFLAG:LeakedCustomerData2] [] [ListAzureResources]
TABLE VI: Agent state at Episode 20, Step 957. The node GitHubProject is discovered (discovered) through a remote exploit from a previously owned node, demonstrating initial expansion of the visible attack surface.
id status properties local_attacks remote_attacks
client owned [] [SearchEdgeHistory] []
Website owned [MySql, Ubuntu, nginx/1.10.3] [CredScanBashHistory] [ScanPageSource, ScanPageContent]
Website[user=monitor] owned [MySql, Ubuntu, nginx/1.10.3] [CredScan-HomeDirectory] []
AzureStorage owned [CTFFLAG:LeakedCustomerData] [] [AccessDataWithSASToken]
AzureResourceManager owned [CTFFLAG:LeakedCustomerData2] [] [ListAzureResources]
GitHubProject discovered [CredScanGitHistory]
TABLE VII: Agent state at Episode 20, Step 958. The agent reveals Website.Directory via directory traversal, highlighting application-layer reconnaissance and continued exploration beyond direct exploitation.
id status properties local_attacks remote_attacks
client owned [] [SearchEdgeHistory] []
Website owned [MySql, Ubuntu, nginx/1.10.3] [CredScanBashHistory] [ScanPageSource, ScanPageContent]
Website[user=monitor] owned [MySql, Ubuntu, nginx/1.10.3] [CredScan-HomeDirectory] []
AzureStorage owned [CTFFLAG:LeakedCustomerData] [] [AccessDataWithSASToken]
AzureResourceManager owned [CTFFLAG:LeakedCustomerData2] [] [ListAzureResources]
GitHubProject discovered [CredScanGitHistory]
Website.Directory discovered [NavigateWebDirectory, NavigateWebDirectoryFurther]
TABLE VIII: Agent state at Episode 20, Step 959. The node Sharepoint is discovered through sequential reconnaissance actions, evidencing the agent’s multi-step lateral movement and expanding visibility into deeper network segments.
id status properties local_attacks remote_attacks
client owned [] [SearchEdgeHistory] []
Website owned [MySql, Ubuntu, nginx/1.10.3] [CredScanBashHistory] [ScanPageSource, ScanPageContent]
Website[user=monitor] owned [MySql, Ubuntu, nginx/1.10.3] [CredScan-HomeDirectory] []
AzureStorage owned [CTFFLAG:LeakedCustomerData] [] [AccessDataWithSASToken]
AzureResourceManager owned [CTFFLAG:LeakedCustomerData2] [] [ListAzureResources]
GitHubProject discovered [CredScanGitHistory]
Website.Directory discovered [NavigateWebDirectory, NavigateWebDirectoryFurther]
Sharepoint discovered [ScanSharepointParentDirectory]
Intermediate Exploration Trajectory of the RL Attacker Under Partial Observability
To illustrate the agent’s behaviour in CTF environment under partial observability, Tables VVIII present four consecutive snapshots of its internal state at Episode 20, Steps 956 to 959. Each row corresponds to a node known to the agent at that step, and the columns represent key aspects of the agent’s perception and action space:
  • id: Unique identifier of a node (e.g., Website, GitHubProject).
  • status: Agent’s access level to the node (owned, discovered, or unknown).
  • properties: Node attributes such as OS, services, or flags (e.g., Ubuntu, nginx, CTFFLAG).
  • local_attacks: Exploits executable from within the node (e.g., CredScanBashHistory, SearchEdgeHistory).
  • remote_attacks: Actions targeting other nodes from the current node (e.g., ScanPageSource, NavigateWebDirectory).
These tables illustrate how the agent progressively uncovers the network topology through a sequence of exploratory actions involving local exploits and remote scans.
At Step 956 (Table V), the agent has compromised five nodes, Website, Website[user=monitor], AzureStorage, AzureResourceManager, and client, each marked as owned. The table lists the local and remote actions currently available from these nodes, including CredScanBashHistory, SearchEdgeHistory, ScanPageSource and others. Now, the agent will execute additional actions from the owned nodes to discover new nodes or compromise further targets, gradually expanding its visibility into the environment.
At Step 957 (Table VI), the agent expands its knowledge of the environment by discovering a new node, GitHubProject, which is now marked as discovered. This node has not yet been compromised, but its appearance indicates that it has been identified as a reachable target through action executed from one of the owned nodes.
At Step 958 (Table VII), the agent continues its exploration by discovering a new node, Website.Directory, which is marked as discovered. This node is revealed through actions executed from an owned node. While the node has not been compromised, its discovery indicates that the agent is actively probing the structure of the environment to uncover additional entry points.
At Step 959 (Table VIII), the agent continues its exploration by discovering a new node, Sharepoint, which is now marked as discovered. Although the node remains uncompromised, its appearance in the state table indicates that the agent is systematically expanding its reach by chaining reconnaissance actions across accessible paths. At this point, the agent has visibility over eight nodes in total, five marked as owned and three as discovered, demonstrating a steady progression in environmental awareness and exploration capabilities.
These four steps represent an intermediate segment within the agent’s broader training trajectory. Over the course of training, the agent incrementally learns to navigate the environment under partial observability by discovering new nodes, exploiting available actions, and updating its internal state representation. This process, combining exploration and exploitation, enables the agent to gradually expand its understanding of the network and identify strategic paths for lateral movement. Examining such intermediate snapshots highlights the underlying mechanisms by which the agent refines and adapts its attack policy over time.